• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ramesh

    Log entry time 08:08:57 on March20,2005

    Entry number 141997

    keyword=Shift Summary

    00:00 Started shift with the Run 2813 from previous shift.


    All shift production runs in

    00:15-00: 00:50 Run 2814 (146k)

    00:47 left Q1 tripped due to instantaneously zero level of Helium flow. In a minute the flow went to normal.

    Could not reset Q1 remotely. In the meantime MCC called that the beam will be off due to Hall B problem. Called Peter.Concluded to request controlled access to reset Q1 manually from the hall.

    02:15 came out of hall resetting Q1 (went into the hall with Peter).
    02:28 went to beam permit
    02:32 Q1 back to normal, asked for beam, MCC said it might take 5 or 7 minutes more since the Hall B problem still persits (which was the cause of taking beam away).

    02:40 beam back.

    02:40-03:27 Run 2815 (200k)
    03:27-03:55 Run 2816 (122k)
    03:55 Q1 tripped again. Helium level drops off to zero and comes back to normal condition in a matter of minute causing the Q1 trip.

    Run 2817 junk

    04:32 Requested 40uA beam

    04:34-05:21 Run 2818 (200k)
    05:22-05:29 Run 2819 (30k)

    05:30 Q1 tripped (third time), could not reset from counting room,requsted escorted access

    06:09 Q1 back after mannual reset in hall, requested 40uA beam.

    06:10-06:54 Run 2820(200k)
    06:54-07:37 Run 2821(200k)
    07:37- Run 1822 passed to day shift.